# Ruelle-Pollicott resonances of diffusive U(1)-invariant qubit circuits

### v1

@@ Abstract
I Introduction
One of the reasons why physics is so successful in describing Nature is because oftentimes simple descriptions suffice. This is so because even simple rules, given for instance by a Hamiltonian, can lead to complicated behavior. How and which simple rules lead to complex behavior is studied within a theory of dynamical systems. Some systems (rules) lead to predictable regular dynamics, the extreme case being integrable systems, while other lead to complex behavior, the extreme case being chaotic systems for which predicting future for a specific initial condition becomes impossible due to exponential complexity growth Ott (2002).
In many-body quantum systems complexity can grow exponentially also with the system size. Such exponentially growing Hilbert space size is what allows for outperformance of quantum devices over classical, and indeed, one of premier usages of present day digital quantum computers and analog simulators is to simulate quantum many-body systems Daley et al. (2022). An interesting question is what new effects can exponential complexity due to chaos and the one from many-body nature bring? To begin with, it is not entirely clear how to define or quantify quantum chaos in many body quantum systems. One can try to use theoretically appealing concepts, like dynamical entropies, which are difficult to use in practice, or one can use a concept that is easy to study numerically, for instance various spectral quantifiers Haake (2010). Both have a drawback that in a many-body system they are hardly observable quantities.
One can, however, ditch theoretical elegance for practical relevance and try to directly study measurable objects. One such are correlation functions
@@ I Introduction
[…]
@@
While many different methods have been proposed and employed to study transport in many-body quantum systems, for a review see Ref. Bertini et al. (2021), here we propose and test a new one that is based on dynamics of operators. It directly works in the thermodynamic limit, i.e., in infinitely-sized systems, with the only approximation being truncating dynamics to operators with a finite local support size. It builds on a decades old idea of Ruelle-Pollicott (RP) resonances Ruelle (1986); Pollicott (1985); Gaspard (1998) that we upgrade and adapt to translationally invariant many-body quantum circuits so that it allows for an easy identification of the transport dynamical exponent, and, for instance, calculation of the diffusion constant.
@@
Figure 1: Diagram of the RP resonance spectrum in systems without (a) and with (b) one conserved quantity. Labels in figure (b) show which parts of the spectrum are responsible for which physical behavior of the system. In the thermodynamic limit, $k$ is continuous and periodic, we normalize it in a way that $k\in[-\pi,\pi)$. We show only positive $k$, since $|\lambda(k)|=\left\lvert\lambda(-k)\right\rvert$, for details see Sec. II.2. The shaded region in figure (b) denotes the location of conjectured continuums of RP resonances, which we cannot observe numerically with used methods.
I.1 Ruelle-Pollicott resonances
The idea of RP resonances is to get the decay rate $\nu$ of exponential asymptotic decay of correlation functions,
$C_{AB}(t)\sim{\rm e}^{-\nu t}=\lambda^{t},$ (2)
where we introduced $\lambda\coloneq{\rm e}^{-\nu}$. While specific observables can decay with their own rate, there is the slowest decay rate $\nu$, the inverse of which is nothing but the system’s relaxation time. It was shown in Refs Pollicott (1985); Ruelle (1986) that in a class of strongly chaotic classical systems $\lambda$ does not dependent on the chosen observable, but is rather an intrinsic property of the system. In such cases, $\lambda$ is referred to as the (leading) RP resonance of the system. While one could get $\lambda$ by simply calculating the appropriate correlation function, the idea is to get $\lambda$ directly from the generator of dynamics (after all, $\lambda$ is an inherent property of the system, not observables).
With classical systems in mind, the starting point is the propagator $\mathcal{U}$, either the Frobenius-Perron propagator of densities, or the Koopman propagator of observables Gaspard (1998); Braun (2001). $\mathcal{U}$ is unitary in the appropriate space (e.g., in the space of $L^{2}$ functions), its eigenvalues lying on the unit circle. One can, however, find poles in the analytic continuation of the resolvent $\mathcal{R}(z)\coloneq 1/(z-\mathcal{U})$ inside the unit circle. It was shown in specific systems Hasegawa and Saphir (1992) that they correspond exactly to RP resonances, the largest being the mentioned $\lambda$.
Performing the analytic continuation for generic systems is not feasible. A more practical approach, which can be carried out either analytically or numerically, is introducing some kind of dissipation or coarse-graining (i.e., non-unitarity) in our system. This leads to a non-unitary propagator $\mathcal{U}(\varepsilon)$, where $\varepsilon$ measures the strength of the non-unitarity. The eigenvalues of $\mathcal{U}(\varepsilon)$ now lie inside the unit circle and, analogous to the RP resonances in the unitary case, govern the decay of generic correlation functions. It is conjectured that in the unitary limit $\varepsilon\to 0$, contrary to a naive expectation that the eigenvalues uniformly converge to the unit circle, this does not happen – they converge to a smaller radius inside the unit circle, which corresponds to an RP resonance. This procedure was used in classical systems, with the non-unitarity being either white noise Gaspard et al. (1995) or a phase space coarse-graining procedure Weber et al. (2000).
Less is known about RP resonances of quantum systems. Initially, they have been studied in the semi-classical regime Pance et al. (2000); Manderfeld (2003); García-Mata et al. (2018). More recently, weak Lindblad dissipation (an analog of noise in classical systems) has been suggested Mori (2024) as a method to extract $\lambda$ via the Liouvillian gap in the limit of infinite systems size and zero noise, with the results for few specific solvable models Jacoby et al. (2024); Zhang et al. (2024); Yoshimura and Sá (2025) agreeing with $\lambda$ calculated directly.
We shall use a different path by considering the Heisenberg propagator of observables $\mathcal{U}(A)\coloneq U^{\dagger}AU$, where $U$ is the unitary propagator, truncating $\mathcal{U}$ to an appropriate operator subspace. In this setting, the RP resonance is expected to correspond to the leading eigenvalue of $\mathcal{U}$. The method was first used on translationally invariant local operators in Ref. Prosen (2002) (which can be though of us akin to a phase space coarse-graining), in the kicked Ising model Prosen (2002, 2004, 2007), and more recently followed by additionally resolving a quasi-momentum Žnidarič (2024). We shall use this quasi-momentum resolved truncated propagator. The truncated propagator can be very handy not just in studies of chaotic systems, but more broadly, for instance, identifying unknown constants of motion in cellular automata Klobas ; Sharipov et al. (2025), or discovering that all homogeneous U(1) conserving quantum circuits are integrable Žnidaric et al. (2024); Žnidarič (2025a).
We note that the recently much discussed Krylov space method can be viewed as a particular way of truncation (in powers of time in the Taylor expansion, i.e., nested commutators) and can thus be related to RP resonances Teretenkov et al. (2025); Loizeau et al. (2025).
I.2 Summary of results
@@
In this work we use the momentum-resolved truncated propagator approach to study RP resonances in systems with one conserved quantity. We focus on many-body quantum circuits with a local U(1) conserved quantity and local gates. Our generic conclusions are demonstrated on magnetization-conserving circuits with 3-site interaction and 3-site translational invariance. Such systems with their associated magnetization transport are of direct experimental interest for NISQ machines Keenan et al. (2023); Maruyoshi et al. (2023); Rosenberg et al. (2024), and are the simplest theoretical setting with a conserved quantity  [^1] Implementation of the truncated propagator method is simpler for quantum circuits compared to Hamiltonian systems due to a sharp light-cone. In addition, one has a 1-local conserved magnetization density instead of a 2-local energy density..
@@
As mentioned, we use the truncated operator propagator Prosen (2002) in its momentum-resolved form Žnidarič (2024), which will allow us to extract transport properties from the momentum dependence of RP resonances. Some work has been done on the connection between RP resonances and diffusion in classical systems Gaspard (1992); Khodas et al. (2000), and conserved quantities have been briefly mentioned in previous Lindblad weak dissipation works Mori (2024); Yoshimura and Sá (2025), as well as for a truncated propagator of the kicked Ising model in a prethermalized regime (or the Hamiltonian limit) in a recent preprint Žnidarič (2025b).
@@
We provide the first comprehensive study of the $k$-dependence of RP resonances, showing that (i) the quasi-momentum dependence of the leading eigenvalue of the truncated propagator and thus the leading RP resonance $\lambda_{1}(k)$ is in general non-trivial (Fig. 1a), even in a system without any conserved quantities where $|\lambda_{1}(k)|$ is gapped away from $1$. For instance, the decay need not be always the fastest at $k=0$, i.e., for translationally invariant observables. (ii) In systems with conserved magnetization one expectedly has $\lambda_{1}(0)=1$ with the corresponding eigenvector being the conserved magnetization, and, more interestingly, from $\lambda_{1}(k)$ for small $k$ one can extract the transport dynamical exponent, and, specifically in the case of diffusion, the value of the diffusion constant due to $|\lambda_{1}(k)|={\rm e}^{-Dk^{2}}$ (Fig. 1b). At large $k$, the diffusive nature of the leading RP resonance ends due to an eigenvalue crossing at $k_{c}$, resulting in an exponential decay of generic observables with sufficiently large quasi-momentum $k$. (iii) We discuss and conjecture a RP continuum below $|\lambda_{1}(k)|$ (Fig. 1b), that are related to non-exponential decay of correlation functions and corrections to diffusive transport.
II Truncated quasi-momentum-dependent propagator of extensive observables
@@
In this section, we describe the quasi-momentum-dependent propagator of local observables, first introduced for $k=0$ in Ref. Prosen (2007) and extended to any $k$ in Ref. Žnidarič (2024). We are considering qubit circuits with propagator for $1$ unit of time $U$ that has translational invariance by $s$ sites, and that can, due to a sharp light-cone, spread local operators supported on $r$ consecutive sites by at most $\delta r$ sites at each edge. For instance, a brickwall circuit composed of all the same gates (Fig. 2a) is invariant under translation by $s=2$ sites and has $\delta r=2$, while a 3-layered brickwall circuit with 3-qubit gates (Fig.3a) has $s=3$ and $\delta r=6$. Previous implementations Prosen (2007); Žnidarič (2024) were for the kicked Ising model that has $s=1$ and $\delta r=1$, or for 2-layer brickwall Žnidarič (2025a); Žnidaric et al. (2024) circuits that have $s=2$ and $\delta r=2$. We generalize it to work for translationally invariant quantum circuits with any geometry, i.e., any $s$ and $\delta r$.
We first express the Heisenberg propagator of observables
@@
in a basis of an infinite system where the quasi-momentum $k$ is a good quantum number. In quantum circuits, the local support $r$ of the density of observable $A$ can increase only by at most $\delta r$ sites at each end, to $r+2\delta r$. See for example Fig. 2 for the brickwall circuit where $\delta r=2$. We then truncate (i.e., project) the observable $\mathcal{U}(A)$ back to observables with densities supported on at most $r$ sites, thus obtaining a non-unitarity truncated propagator $\mathcal{U}^{(r)}$. In the limit of the non-unitarity going to zero, that is $r\to\infty$, we expect to extract RP resonances as the “frozen” eigenvalues of the truncated propagator $\mathcal{U}^{(r)}$.
II.1 Brief description
@@
where $\mathcal{S}$ is the 1-site translation super-operator to the right (i.e., $\mathcal{S}(a\otimes\mathbbm{1})=\mathbbm{1}\otimes a$), $a$ is a local operator and the sum runs over all $j\in\{0,\dots,N/s-1\}$ possible translations by $s$ sites. In a finite system, $k$ takes values $\frac{2\pi s}{N}j$, however, at the end we will be interested in the thermodynamic limit $N\to\infty$  [^2] Formally, RP resonances can only exist in systems with infinitely dimensional Hilbert spaces. In finite systems, correlation functions always converge to a nonzero constant., where $k\in[0,2\pi)$. One can check that $A$ really is an eigenvector of the translation super-operator, $\mathcal{S}(A)={\rm e}^{{\rm i}k}A$. Such $A$ will be referred to as an extensive observable.
We wish to work with local operators $a$ as proxies for extensive observables $A$. It is thus crucial to restrict ourselves to a space of local operators in which the correspondence between local operators and extensive observables is one-to-one. Namely, in Eq. (4), different $a$ can result in the same $A$. For example, both $a=\sigma^{z}_{1}$ and $a=\frac{1}{2}\left(\sigma^{z}_{1}+\sigma^{z}_{1+s}\right)$ represent the same $A$. Here $\sigma_{j}^{\alpha}$ denotes the $\alpha$ Pauli matrix acting on site $j$. To avoid that and have a one-to-one correspondence, such translations of components of $a$ by $s$ sites must be prohibited. This can be done in various ways, we choose to write the extensive observable in the following form
@@
where $\mathcal{S}^{sj+m}(a_{m})$ can be thought of as operators starting on sites $sj+m$ (i.e., on all sites $i$, where $i\ (\mathrm{mod})\ s=m$). In other words, due to translational invariance by $s$ sites we have a local unit call of $s$ sites, so that the sum over $m$ runs over sites within a unit cell, while $j$ runs over all unit cells. In this formulation it is now easy to see that we will have a one-to-one correspondence between the set of $\{a_{m}\}_{m=0}^{s-1}$ and extensive observable $A$ if we in addition enforce that $a_{m}$ do not contain components acting trivially on the first site. Explicitly, $a_{m}$ must be in the space spanned by the set (basis) $\mathcal{P}^{(r)}$ of Pauli strings without identities acting on the first site:
$\mathcal{p}_{0}$ :-$\coloneq\{\mathbbm{1}\}\cup\mathcal{p},\qquad\text{where}\qquad\mathcal{p}\coloneq\{\sigma^{x},\sigma^{y},\sigma^{z}\},:$ (6)
$\mathcal{P}^{(r)}$ :-$\coloneq\left\{p_{1}\otimes p_{2}\otimes\cdots\otimes p_{r}\ |\ p_{1}\in\mathcal{p}\ \land\ p_{j\neq 1}\in\mathcal{p}_{0}\right\}.:$
@@
where $b\in\mathcal{P}^{(r)}$ for some $r$. Moreover, $B_{k}^{(m,b)}$ for all $k,m$ and $b\in\mathcal{P}^{(r)}$ are an orthonormal basis of the space of extensive observable w.r.t. the extensive Hilbert-Schmidt inner product $\left\langle\!\left\langle A,B\right\rangle\!\right\rangle\coloneq\frac{s}{N2^{N}}\mathrm{tr}\,A^{\dagger}B$. The number of basis elements of support $r$ (or less) is $\frac{3s}{4}4^{r}$. Importantly, the identity operator $\mathbbm{1}$ is excluded from this basis. This is convenient, since it is always a conserved quantity in unitary dynamics and is, therefore, not of interest in the present application. For details and derivations see Appendix A.1.
@@
Refer to caption(a)                                                 Refer to caption (b)                                                Refer to caption(c)
Figure 2: Circuits with no conserved quantities. (a) Diagram of operator spreading in a brickwall circuit with the same gate $V$ acting between all nearest neighbors (the shown circuit has $s=2$, $\delta r=2$). (b) The leading eigenvalue of the propagator truncated to the space of extensive observables with support $r$ for two cases of a circuit, with $V$ in each case chosen randomly according to the unitary Haar measure. (c) Infinite temperature autocorrelation functions of extensive observables $A$ (4) in the circuit denoted as Case i in (b). The local density $a$ is either $\sigma^{z}$, or chosen randomly according to the Gaussian unitary ensemble (GUE) Mehta (2004). Dashed lines show the RP prediction $\propto\left\lvert\lambda_{1}\right\rvert^{t}$, while full curves are exact calculation in a circuit with $N=30$ qubits; the $\sigma^{z},k=0$ case is multiplied by $15$ for better presentation.
II.2 Truncation and Ruelle-Pollicott resonances
The Heisenberg propagator in Eq. (9) evaluated in the entire (infinitely dimensional) Hilbert space is a unitary operator. We wish to introduce a non-unitary by restricting ourselves to some subspace. There are different options, one of the most natural ones is restricting ourselves to the space of support $r$ defined in Eq. (6). Projections of the Heisenberg propagator to this space results in the truncated propagator of extensive observables $\mathcal{U}^{(r)}_{k}$ with matrix elements

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
One of the reasons why physics is so successful in describing nature is because oftentimes simple descriptions suffice. This is so since even simple rules, given for instance by a Hamiltonian, can lead to complicated behavior. How and which simple rules lead to complex behavior is studied within the theory of dynamical systems. Some systems (rules) lead to predictable regular dynamics, the extreme case being integrable systems, while others lead to complex behavior, the extreme case being chaotic systems for which predicting the future for a specific initial condition becomes impossible due to exponential complexity growth [31].
In many-body quantum systems, complexity can grow exponentially also with the system size. Such exponentially growing Hilbert space size is what allows for outperformance of quantum devices over classical, and indeed, one of the premier uses of present day digital quantum computers and analog simulators is to simulate quantum many-body systems [4]. An interesting question is: What new effects can arise from the combination of exponential complexity due to chaos and that from the many-body nature? To begin with, it is not entirely clear how to define or quantify quantum chaos in many-body quantum systems. One can try to use theoretically appealing concepts, like dynamical entropies, which are difficult to use in practice, or one can use a concept that is easy to study numerically, for instance one of the various spectral quantifiers [12]. Both have a drawback that in a many-body system they are hardly observable quantities.
One can, however, abandon theoretical elegance for practical relevance and try to directly study measurable objects. An example is the correlation function
@@ 1 Introduction
where $\left\langle\bullet\right\rangle$ denotes the average over an appropriate ensemble. A defining property of mixing systems is that correlation functions decay to zero, while for chaotic systems we usually expect more, namely, that $C_{AB}(t)$ decay exponentially at long enough times, $C_{AB}(t)\sim{\rm e}^{-\nu t}$. Time-dependent correlation functions specifically allow us to study dynamics, and through linear response also get access to non-equilibrium properties. A prime example is the transport of globally conserved quantities, like energy, charge, or magnetization.
@@
While many different methods have been proposed and employed to study transport in many-body quantum systems, for a review see Ref. [1], here we propose and test a new one that is based on the dynamics of operators. It directly works in the thermodynamic limit, i.e., in infinite-sized systems, with the only approximation being truncating the dynamics to operators with a finite local support size. It builds on a decades-old idea of Ruelle-Pollicott (RP) resonances [40, 34, 11] that we upgrade and adapt to translationally invariant many-body quantum circuits, so that it allows for easy an calculation of the transport dynamical exponent and phenomenological constants, for instance, the diffusion constant.
1.1 Ruelle-Pollicott resonances
The idea of RP resonances is to get the decay rate $\nu$ of the asymptotic exponential decay of correlation functions,
$C_{AB}(t)\sim{\rm e}^{-\nu t}=\lambda^{t},$ (2)
where we introduced $\lambda\coloneq{\rm e}^{-\nu}$. While specific observables can decay with their own rate, there is the slowest decay rate $\nu$, whose inverse is the system’s relaxation time. It was shown in Refs. [34, 40] that in a class of strongly chaotic classical systems $\lambda$ does not depend on the chosen observable, but is rather an intrinsic property of the system. In such cases, $\lambda$ is referred to as the (leading) RP resonance of the system. While one could get $\lambda$ by simply calculating the appropriate correlation function, the idea is to get $\lambda$ directly from the generator of dynamics (after all, $\lambda$ is an inherent property of the system, not of the observables).
With classical systems in mind, the starting point is the propagator $\mathcal{U}$, either the Frobenius-Perron propagator of densities, or the Koopman propagator of observables [11, 2]. $\mathcal{U}$ is unitary in the appropriate space (e.g., in the space of $L^{2}$ functions), its eigenvalues lying on the unit circle. One can, however, find poles in the analytic continuation of the resolvent $\mathcal{R}(z)\coloneq 1/(z-\mathcal{U})$ inside the unit circle. It was shown in specific systems [13] that they correspond exactly to RP resonances, the largest being the mentioned $\lambda$.
Performing the analytic continuation for generic systems is not feasible. A more practical approach, which can be carried out either analytically or numerically, is introducing some kind of dissipation or coarse-graining (i.e., a non-unitarity) in our system. This leads to a non-unitary propagator $\mathcal{U}(\varepsilon)$, where $\varepsilon$ measures the strength of the non-unitarity. The eigenvalues of $\mathcal{U}(\varepsilon)$ now lie inside the unit circle and, analogous to the RP resonances in the unitary case, govern the decay of generic correlation functions. It is conjectured that in the unitary limit $\varepsilon\to 0$ the naive expectation that the eigenvalues uniformly converge to the unit circle is not correct – they converge to a smaller radius inside the unit circle, which corresponds to an RP resonance. This procedure was used in classical systems, with the non-unitarity being either white noise [9] or a phase space coarse-graining procedure [46].
Less is known about RP resonances of quantum systems. Initially, they have been studied in the semi-classical regime [33, 23, 8]. More recently, weak Lindblad dissipation (an analog of noise in classical systems) has been suggested [29] as a method of extracting $\lambda$ via the Liouvillian gap in the limit of infinite system size and zero noise, with the results for a few specific solvable models [14, 49, 48] agreeing with $\lambda$ calculated directly.
We shall use a different path by considering the Heisenberg propagator of observables $\mathcal{U}(A)\coloneq U^{\dagger}AU$, where $U$ is the unitary propagator, truncating $\mathcal{U}$ to an appropriate operator subspace. In this setting, the RP resonance is expected to correspond to the leading eigenvalue of $\mathcal{U}$. The method was first used on translationally invariant local operators in Ref. [35] (which can be thought of as akin to a phase space coarse-graining) in the kicked Ising model [35, 36, 37], and more recently followed by additionally resolving the quasi-momentum [53]. We shall use this quasi-momentum-resolved truncated propagator. The truncated propagator can be very handy not just in studies of chaotic systems, but more broadly, for instance, in identifying unknown constants of motion in cellular automata [18, 41], or discovering that all homogeneous U(1) conserving quantum circuits are integrable [52, 50].
We note that the recently much discussed Krylov space method can be viewed as a particular way of truncation (in powers of time in the Taylor expansion, i.e., nested commutators) and can thus be related to RP resonances [42, 22].
1.2 Summary of results
@@
Figure 1: Diagram of the $k$-dependent RP resonance spectrum in homogeneous systems without (a), and with one conserved quantity (b). Labels in figure (b) show which parts of the spectrum are responsible for which physical behavior of the system. The shaded region in figure (b) denotes the location of conjectured continuums of RP resonances, which we cannot observe numerically with the methods used. We show only positive $k$ since $|\lambda(k)|=\left\lvert\lambda(-k)\right\rvert$, for details see Sec. $2.$
@@
In this work we study RP resonances of systems with one conserved quantity. We focus on many-body quantum circuits with a local U(1) conserved quantity and local gates. Such systems, with their associated transport, are of direct experimental interest for NISQ machines [16, 24, 39], and are the simplest theoretical setting with a conserved quantity [^1] Implementation of the truncated propagator method is simpler for quantum circuits compared to Hamiltonian systems due to a sharp light-cone. In addition, one has conserved magnetization with a 1-local density instead of conserved energy with a 2-local density.. Our generic conclusions are demonstrated on magnetization-conserving circuits with 3-site interaction and 3-site translational invariance.
@@
As mentioned, we use the truncated operator propagator [35] in its quasi-momentum-resolved form [53], which allows us to extract transport properties from the momentum dependence of RP resonances. Some work has been done on the connection between RP resonances and diffusion in classical systems [10, 17], and conserved quantities have been briefly mentioned in previous Lindblad weak dissipation works [29, 48], as well as for the truncated propagator of the kicked Ising model in a prethermalized regime (or the Hamiltonian limit) in a recent preprint [51]. We also note that a somewhat related idea of dropping operators with many non-identity Pauli operators has been used in Ref. [38], where the authors introduce a numerical Lindblad evolution method that aims to reach the long-time hydrodynamic behavior.
@@
We provide the first comprehensive study of the $k$-dependence of RP resonances, showing that (i) the quasi-momentum dependence of the leading eigenvalue of the truncated propagator and thus the leading RP resonance $\lambda_{1}(k)$ is in general non-trivial (Fig. 1a), even in a system without any conserved quantities where $|\lambda_{1}(k)|$ is gapped away from $1$. For instance, the decay need not be always the slowest at $k=0$, i.e., for translationally invariant observables. (ii) In systems with conserved magnetization, one expectedly has $\lambda_{1}(0)=1$ with the corresponding eigenvector being the conserved magnetization, and, more interestingly, from $\lambda_{1}(k)$ for small $k$ one can extract the transport dynamical exponent, and, specifically in the case of diffusion, the value of the diffusion constant from $|\lambda_{1}(k)|={\rm e}^{-Dk^{2}}$ (Fig. 1b). At large $k$, the diffusive nature of the leading RP resonance ends due to an eigenvalue crossing at $k_{c}$, resulting in an exponential decay of generic observables with sufficiently large quasi-momentum $k$. (iii) We discuss and conjecture an RP continuum below $|\lambda_{1}(k)|$ (Fig. 1b) that is related to non-exponential decay of correlation functions and corrections to diffusive transport.
2 Truncated quasi-momentum-dependent propagator of extensive observables
@@
In this section, we describe the quasi-momentum-dependent propagator of local observables, first introduced for $k=0$ in Ref. [35] and extended to any $k$ in Ref. [53]. We are considering qubit circuits with propagators for $1$ unit of time $U$ that have translational invariance by $s$ sites, and that can, due to a sharp light-cone, spread local operators supported on $r$ consecutive sites by at most $\delta r$ sites at each edge. For instance, a brickwall circuit composed of all the same gates (Fig. 2a) is invariant under translation by $s=2$ sites and has $\delta r=2$, while a 3-layered brickwall circuit with 3-qubit gates (Fig. 3a) has $s=3$ and $\delta r=6$. Previous implementations [35, 36, 37, 53] were for the kicked Ising model that has $s=1$ and $\delta r=1$, or for the (2-layer) brickwall circuits [50, 52]. We generalize it to work for translationally invariant quantum circuits with any geometry, i.e., any $s$ and $\delta r$.
First, we express the Heisenberg propagator of observables
@@
in an operator basis of an infinite system where the quasi-momentum $k$ is a good quantum number. We then truncate (i.e., project) the observable $\mathcal{U}(A)$ back to observables with densities supported on at most $r$ consecutive sites, thus obtaining a non-unitary truncated propagator $\mathcal{U}^{(r)}$. In the limit of the non-unitarity going to zero, that is $r\to\infty$, we expect to extract RP resonances as the “frozen” eigenvalues of the truncated propagator $\mathcal{U}^{(r)}$.
2.1 Brief description
@@
where $\mathcal{S}$ is the 1-site translation super-operator to the right (i.e., $\mathcal{S}(a\otimes\mathbbm{1})=\mathbbm{1}\otimes a$), $a$ is a local operator and the sum runs over all $j\in\{0,\dots,N/s-1\}$ possible translations by $s$ sites. In a finite system, $k$ takes values $\frac{2\pi s}{N}j$, however, at the end we will be interested in the thermodynamic limit $N\to\infty$ [^2] Formally, RP resonances can only exist in systems with infinite-dimensional Hilbert spaces. In finite systems, correlation functions always converge to a nonzero constant., where we can take $k\in(-\pi,\pi]$. One can check that $A$ really is an eigenvector of the translation super-operator, $\mathcal{S}(A)={\rm e}^{{\rm i}k}A$. Such $A$ will be referred to as an extensive observable.
We wish to work with local operators $a$ as proxies for extensive observables $A$. It is thus crucial to restrict ourselves to a space of local operators in which the correspondence between local operators and extensive observables is one-to-one. Namely, in Eq. (4), different $a$ can result in the same $A$. For example, both $a=\sigma^{z}_{1}$ and $a=\frac{1}{2}\left(\sigma^{z}_{1}+\sigma^{z}_{1+s}\right)$ represent the same $A$. Here $\sigma_{j}^{\alpha}$ denotes the $\alpha$ Pauli matrix acting on site $j$. To have a one-to-one correspondence, such translations of components of $a$ by $s$ sites must be prohibited. This can be done in various ways, we choose to write the extensive observable in the following form
@@
where $\mathcal{S}^{sj+m}(a_{m})$ can be thought of as operators starting on sites $sj+m$ (i.e., on all sites $i$, where i(mod)s=mi\ ($\mathrm${mod})\ s=m). In other words, due to translational invariance by $s$ sites we have a local unit cell of $s$ sites, so that the sum over $m$ runs over sites within a unit cell, while $j$ runs over all unit cells. In this formulation it is now easy to see that we have a one-to-one correspondence between the set of $\{a_{m}\}_{m=0}^{s-1}$ and the extensive observable $A$ if we enforce that $a_{m}$ do not contain components acting trivially on the first site. Explicitly, $a_{m}$ must be in the space spanned by the set (basis) $\mathcal{P}^{(r)}$ of Pauli strings without identities acting on the first site:
$\mathcal{p}_{0}$ $\coloneq\{\mathbbm{1}\}\cup\mathcal{p},\qquad\text{where}\qquad\mathcal{p}\coloneq\{\sigma^{x},\sigma^{y},\sigma^{z}\},$ (6)
$\mathcal{P}^{(r)}$ $\coloneq\left\{p_{1}\otimes p_{2}\otimes\cdots\otimes p_{r}\ |\ p_{1}\in\mathcal{p}\ \land\ p_{j\neq 1}\in\mathcal{p}_{0}\right\}.$
@@
where $b\in\mathcal{P}^{(r)}$ for some $r$. Moreover, $B_{k}^{(m,b)}$ for all $k,m$ and $b\in\mathcal{P}^{(r)}$ are an orthonormal basis of the space of extensive observables w.r.t. the extensive Hilbert-Schmidt inner product, $\left\langle\!\left\langle A,B\right\rangle\!\right\rangle\coloneq\frac{s}{N2^{N}}\mathrm{tr}\,A^{\dagger}B$. The number of basis elements of support $r$ (or less) is $\frac{3s}{4}4^{r}$. Importantly, the identity operator $\mathbbm{1}$ is excluded from this basis. This is convenient, since it is always a conserved quantity in unitary dynamics and is, therefore, not of interest in the present application. For details and derivations regarding the space of extensive observables see Appendix A.1.
@@
2.2 Truncation and Ruelle-Pollicott resonances
The Heisenberg propagator in Eq. (9) evaluated on the entire (infinite-dimensional) Hilbert space is a unitary operator. We wish to introduce a non-unitarity by restricting ourselves to some subspace. There are different options, one of the most natural ones is restricting ourselves to the space of support $r$ defined in Eq. (6). Projection of the Heisenberg propagator to this space results in the truncated propagator of extensive observables $\mathcal{U}^{(r)}_{k}$ with matrix elements